Assume that we want to construct a confidence interval. Do one of the​ following, as​ appropriate: (a) find the critical value t Subscript alpha divided by 2​, ​(b) find the critical value z Subscript alpha divided by 2​, or​ (c) state that neither the normal distribution nor the t distribution applies. The confidence level is 90​%, sigmaequals3773 thousand​ dollars, and the histogram of 64 player salaries​ (in thousands of​ dollars) of football players on a team is as shown.
Answer #1

Solution

Given that,

Standard deviation = = 3773

Sample size =n = 64

At 90% confidence level the z is ,

= 1 - 90% = 1 - 0.90 = 0.10

/ 2 = 0.10 / 2 = 0.05

Z/2 = Z0.05 = 1.645

The critical value z* = 1.645